Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Three students, Mark, Peter, and Wanda, are all working on the same math problem. If their individual probabilities of

Expert replies
by BTGmoderatorDC » Thu Jan 16, 2020 6:24 pm
Three students, Mark, Peter, and Wanda, are all working on the same math problem. If their individual probabilities of success are 1/4, 2/5, and 3/8, respectively, then what is the probability that at least one of the students will get the problem correct?

A. 3/80
B. 9/32
C. 23/32
D. 77/80
E. 39/40



OA C

Source: Princeton Review
Join the discussion
Source: — Problem Solving |

The problem will be solved in three conditions:

1. Anyone among Mark, Peter, and Wanda solves the problem and the other two don't
2. Any two among Mark, Peter, and Wanda solve the problem and the other one doesn't
3. All three Mark, Peter, and Wanda solve the problem

So, the problem will remain unsolved only when all three are not able to solve it. So, a better approach to get to the question is to find out the probability that none could solve the problem and then deduct it from 1.

• Probability of Mark not able to solve the problem = 1 – 1/4 = 3/4;
• Probability of Peter not able to solve the problem = 1 – 2/5 = 3/5;
• Probability of Wanda not able to solve the problem = 1 – 3/8 = 5/8

Thus, the probability that none can solve the problem = (3/4)*(3/5)*(5/8) = 9/32

Thus, the probability that at least one can solve the problem = 1 – 9/32 = 23/32

The correct answer: C

Hope this helps!

-Jay
_________________
Manhattan Review

Locations: Manhattan Review Visakhapatnam | GMAT Prep Warangal | GRE Prep Dilsukhnagar | Begumpet GRE Coaching | and many more...

Schedule your free consultation with an experienced GMAT Prep Advisor! Click here.
Join the discussion

BTGmoderatorDC wrote: ↑
Thu Jan 16, 2020 6:24 pm
Three students, Mark, Peter, and Wanda, are all working on the same math problem. If their individual probabilities of success are 1/4, 2/5, and 3/8, respectively, then what is the probability that at least one of the students will get the problem correct?

A. 3/80
B. 9/32
C. 23/32
D. 77/80
E. 39/40



OA C

Source: Princeton Review
CASE 1
Mark right \(1/4\); wrong \(3/4\)
Peter right \(2/5\); wrong \(3/5\)
Wanda right \(3/8\); wrong \(5/8\)

So, we get following cases:
\(1/4\cdot 3/5\cdot 5/8 + 2/5\cdot 3/4 \cdot 5/8 + 3/4\cdot 3/5\cdot 3/8 + 1/4\cdot 2/5\cdot 5/8 + 1/4\cdot 3/8\cdot 3/5 + 2/5\cdot 3/8\cdot 3/4 + 1/4\cdot 2/5\cdot 3/8\); \(115/160\); \(23/32\). Therefore, __C__

CASE 2
Else all wrong cases, \(3/4\cdot 3/5\cdot 5/8 = 9/32\)
1- wrong cases ; \(1-9/32; 23/32\). Therefore, __C__
Join the discussion

BTGmoderatorDC wrote: ↑
Thu Jan 16, 2020 6:24 pm
Three students, Mark, Peter, and Wanda, are all working on the same math problem. If their individual probabilities of success are 1/4, 2/5, and 3/8, respectively, then what is the probability that at least one of the students will get the problem correct?

A. 3/80
B. 9/32
C. 23/32
D. 77/80
E. 39/40



OA C

Source: Princeton Review
We can use the formula:

1 - P(none get the problem correct) = P(at least one gets the problem correct)

1 - P(none get the problem correct) = 1 - 3/4 x 3/5 x 5/8 = 1 - 9/32 = 23/32

Answer: C

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion

BTGmoderatorDC wrote: ↑
Thu Jan 16, 2020 6:24 pm
Three students, Mark, Peter, and Wanda, are all working on the same math problem. If their individual probabilities of success are 1/4, 2/5, and 3/8, respectively, then what is the probability that at least one of the students will get the problem correct?

A. 3/80
B. 9/32
C. 23/32
D. 77/80
E. 39/40



OA C

Source: Princeton Review
Aside: If P(Mark is correct) = 1/4, then P(Mark is INcorrect) = 3/4
If P(Peter is correct) = 2/5, then P(Peter is INcorrect) = 3/5
If P(Wanda is correct) = 3/8, then P(Wanda is INcorrect) = 5/8
---------------------------

When it comes to probability questions involving at least, it's best to try using the complement.
That is, P(Event A happening) = 1 - P(Event A not happening)

So, we can write: P(at least one student is correct) = 1 - P(NO students are correct)

P(NO students are correct) = P(Mark is incorrect AND Peter is incorrect AND Wanda is incorrect)
= P(Mark is incorrect) x P(Peter is incorrect) x P(Wanda is incorrect)
= 3/4 x 3/5 x 5/8
= 9/32

So, P(at least one student is correct) = 1 - 9/32
= 23/32

Answer: C

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion